OpenStudy (anonymous):

The function f(x)=2x/(x^2+1) is 1.)Even 2.)Odd 3.)Neither even nor odd 4.)The concept of being even or odd does 5.)not apply since this expression does not represent a function 6.)It is both even and odd.

OpenStudy (anonymous):

If the Function is Even then f(-x) = f(x) if the function is odd then f(-x) = -f(x) hope this helps

OpenStudy (anonymous):

so its not even or odd

OpenStudy (anonymous):

choice 3

OpenStudy (anonymous):

Did u Check ..?

OpenStudy (anonymous):

how do i check

OpenStudy (anonymous):

By putting x = -x and Check...i think..it is odd...

OpenStudy (anonymous):

huh im still a little lost

OpenStudy (anonymous):

It is an odd function... f(-x) = -x/x^2 + 1 = -f(x)

OpenStudy (anonymous):

is that how you would solve the problem

OpenStudy (anonymous):

This is the actual way we solve this kind of problems, by replacing x with (-x) and checking whether f(x) is equal to f(x) or -f(x)...

OpenStudy (anonymous):

If f(x) = -f(x) . an odd function if f(x) = +f(x) an even function Hope this makes sense....

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

medal please...